OpenStudy (anonymous):

simplify: 5x squared - 20x squared - 80x squared

OpenStudy (anonymous):

5-20-80

OpenStudy (anonymous):

-95

OpenStudy (anonymous):

\[\sqrt{5x}-\sqrt{20x}-\sqrt{80x}\]

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

thats not one of my answers

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

V(A * X) => V(A) * V(X) Thus: V(5x) - V(20x) - V(80x) = V(5) * V(X) - V(20) * V(x) - V(80) * V(x) so when u substract that you get: -11,18033989 V(X)